OpenStudy (anonymous):

5i/6+3i I know the answer is 1+2i/3 but what is the steps to get there

OpenStudy (smored):

\[\frac{ 5i }{ 6+3i }\] multiply by the denominator's complement (if a+bi, them complement= a-bi) \[\frac{ 5i }{ 6+3i }\ * \frac{ 6-3i }{ 6-3i }\] \[\frac{ 30-(-15) }{ 36-(-9) }\] (i times i equals -1) \[\frac{ 45 }{ 45 } = 1\]

OpenStudy (smored):

oops, made a mistake

OpenStudy (smored):

its \[\frac{ 30i-(-15) }{ 36-(-9) }\]

OpenStudy (smored):

\[\frac{ 30i + 15 }{ 45 } = \frac{ 2+i }{ 3 }\]

OpenStudy (anonymous):

okay I get how you got the 30i +15/45 but how does 30i simplify to 2i+1/3 I keep getting 30i +1/3 I simplified 15/45

OpenStudy (smored):

it is\[\frac{ 30i }{ 45 } + \frac{ 15 }{ 45 }\] 15/45 is 1/3 and 30i/45=2i/3 \[\frac{ 2i }{ 3 } + \frac{ 1 }{ 3 }\] equals to\[\frac{ 2i+1 }{ }\]

OpenStudy (smored):

forgot to put 3 in denominator

OpenStudy (anonymous):

awesome that helps so much

OpenStudy (smored):

no problem!